I need to find the value of v in the following log.
v=log(base 2)(1/4)
If I do as the law says, I come up with -2. But the book is showing 22. Am I doing this wrong or is the books answer wrong.

Answer by khwang(438)

If v=log(base 2)(1/4), by definition of log,
2^v = 4.
Now if v= -2, 2^(-2) = 1/2^2 = 1/4
So your answer is correct.
One way to solve:
v = log(base 2)(1/4) = -log(base 2)4 = -log(base 2) 2^2
= - 2 log(base 2) 2 = -2 * 1 = -2.